One number is three less than twice another. The sum of their squares is 74. Find the numbers.
Answer by stanbon(75887)

One number is three less than twice another. The sum of their squares is 74. Find the numbers.
-----------
1st: x
2nd: 2x-3
-----
Equation:
x^2 + (2x-3)^2 = 74
x^2 + 4x^2 -12x + 9 = 74
---
5x^2 -12x - 65 = 0
---
(x-5)(5x+13) = 0
1st: x = 5
2nd: 2x-3 = 7
===============
Cheers,
stan H.